What is the voltage drop across the 3.0 Q for a parallel circuit containing a 12 V potential and three resistors R1= 3 ohms, R2=2 ohms and R3= 1 ohm


1
Expert's answer
2021-05-09T14:59:35-0400

Given

Potential (V)=12volt

R1=3ohm

R2=2ohm

R3=1ohm

Rnet=R1||R2||R3

Rnet="\frac{6}{11}ohm"

Current

"I=\frac{V}{R}"

"I=\frac{12}{\frac{6}{11}}"

"I=22Amp"

Voltage drop across all three

resistance are parallel circuit

"V_{T}=V_{1}=V_{2}=V_{3}"

R2 corresponding voltage drop

V2=12"volt"